OpenStudy (anonymous):

Find the equation of a straight line that passes through (1, -2) and (-1, 4) Using m= y2 - y1 ------- x2 - x1 I done so far: y = -3x + c, how do i find c? Seeing as I have to find the equation. Any help please?! Thanks

OpenStudy (anonymous):

all my lines are straight.

OpenStudy (anonymous):

the slope is = (4+2)/(-1-1) =-3 Now I choose the point (1,-2) and applying the formula y-y1=m(x-x1) I got y+2 = -3(x-1) y+2=-3x+3 3x+y-10=0 Any question, please ask

OpenStudy (anonymous):

Thank you very much ermunrradtvamg!!!!, this is just what I needed, I was not aware of the second formula and I was starting to get extremely annoyed with this question. You have explained perfectly!

OpenStudy (anonymous):

UW

OpenStudy (anonymous):

Actually sorry one question the end shouldn't it be y = -3x + 1? Seeing as y + 2 = -3x + 3 The 2 will move to the other side and it will become subtract from that to equal: y = -3x + 1